# Self-supervised 3D Point Cloud Completion via Multi-view Adversarial Learning

### v2

@@ IV-A Dataset and Evaluation Metric
Dataset Methods Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
@@
IV-B Implementation Details
The point numbers of $\mathbf{P}_{in}$, $\mathbf{P}_{c}$, and $\mathbf{P}_{out}$ are 2048, 2048 and 16384, respectively. We employ 16 and 24 neighbors for calculating position and curvature encodings, respectively. Initial point radii for the DARE algorithm are set to 0.008 for the plane category and 0.03 for other categories. The loss weights are all set to 1. The initial learning rate is 0.0001, and it decays by 0.5 for every 200 epochs. The training process takes 600 epochs to converge. The experiments are conducted on an NVIDIA RTX 3090 GPU with Intel(R) Xeon(R) CPUs.
IV-C Comparison with State-of-the-Art Methods
IV-C1 Results on synthetic data
@@
We first conduct experiments on the synthetic dataset and compare the results with other state-of-the-art methods, including fully supervised methods (i.e., FoldingNet [41], PCN [15], TopNet [42], AnchorFormer [17]), unpaired methods (i.e., PCL2PCL [25], KTNet [29], LTUPCN [63], USSPA [30]) and self-supervised methods (i.e., ACL-SPC [66], P2C [67]). We train our model in two modes: one is training separately for each category, and the other is training all eight categories together while the adversarial depth maps for each object are from the same category. In contrast, all compared methods are only trained for each category. The quantitative results presented in Table I suggest that our self-supervised method achieves an impressive average CD of 9.08, which outperforms other self-supervised methods and two unpaired approaches (i.e., PCL2PCL and USSPA) by a large margin, and even demonstrating competitive results compared to some fully supervised methods. Regarding the category-specific results, our method significantly outperforms two self-supervised methods across five categories. Specifically, the CD value of MAL-SPC under the lamp category surpasses the results of P2C and ACL-SPC by nearly 5 and 20, respectively. Besides, the CD values of MAL-SPC are lower than those of self-supervised benchmarks by 2-4 under cabinet, couch, and table categories. Furthermore, even with the prior knowledge of complete 3D point clouds, the results of the two unpaired methods, PCL2PCL and USSPA, lag behind our method across multiple categories. In addition, compared to the supervised method FoldingNet, our model achieves a CD difference of less than 3.
As shown in the last two rows of Table I, the results of our method trained separately for each category are generally better than training them together. However, the latter also outperforms other self-supervised methods. Table II shows that our method achieves the best Precision and Coverage across most categories, and outperforms other self-supervised methods in terms of average Precision. Although it falls slightly behind ACL-SPC in Coverage, the difference remains minimal. These results indicate that our method predicts highly complete point clouds while effectively matching local shapes. Table III indicates that our method leads on the PCN and CRN results, with the average CD surpassing the second place by 6.04 and 6.09, respectively. However, the performance of our method and ACL-SPC shows a significant decline on the 3D-EPN dataset. To simulate the human perspective, we set the elevation angle for scanning objects between -30 and 30 degrees when generating our dataset, whereas the partial point clouds in 3D-EPN are scanned from arbitrary angles. This difference in viewpoints introduces a domain gap between the datasets, resulting in the underperformance of our method and ACL-SPC. Note that the performance of P2C increases on the 3D-EPN dataset, indicating that it is more robust to viewpoint variations of partial shape scan.
Moreover, we evaluate the performance of our method in the unpaired case. Specifically, we replace our MAN module with the MLP-based discriminator from USSPA and use the predicted point cloud and an unpaired complete shape as discrimination targets. As shown in Table I, our method achieved favorable results in the unpaired case, with the average CD significantly outperforming PCL2PCL and USSPA. However, the results slightly lag behind our self-supervised approach. We believe that, on the one hand, our PRN effectively learns the common 3D structural features within each object and category, thereby reducing the impact of the unpaired complete point cloud. On the other hand, unlike unpaired methods, our MAN leverages a 2D network as the discriminator. It is well-known that 2D neural networks generally possess stronger learning capabilities compared to 3D MLPs, and our method is able to effectively utilize this advantage.
@@
Figure 4: Qualitative results on the synthetic dataset. (a) Partial points. (b) Ground truth. (c) FoldingNet [41]. (d) PCN [15]. (e) AnchorFormer [17]. (f) KTNet [29]. (g) USSPA [30]. (h) ACLSPC [66]. (i) P2C [67]. (j) Ours.
@@
Figure 4 illustrates a comparison of the visualization results between our method and other benchmarks. We can see from Figure 4 that MAL-SPC can infer a more reasonable global shape while better preserving the incomplete parts. For example, the four legs of the input chair point cloud are all thin cylindrical shapes, while other self-supervised or even supervised methods reconstruct chair legs as thick rectangular prisms, whereas MAL-SPC preserves its original shape well. In addition, it is obvious that the noise points generated by our method are significantly fewer than those produced by other methods. Self-supervised benchmarks and most supervised approaches generate a large number of outlier points outside the edges of objects, whereas our method has almost no outlier points.
@@ IV-C2 Results on real-world scans
In order to validate the robustness and effectiveness of our method in real-world environments, we evaluate its performance on two challenging datasets: ScanNet and KITTI. As demonstrated in Table IV, compared to ACL-SPC and P2C, our MAL-SPC yields the best UCD on all three categories and the best UHD on the car and table categories. Our method only slightly falls behind the second-best approach in terms of UCD on the chair category. Moreover, regarding the results on the car category, MAL-SPC even surpasses the unpaired and supervised methods by a large margin. Figure 5 shows the quantitative results of our model and other state-of-the-art benchmarks. Similar to its superior performance on synthetic data, our method predicts more reasonable shapes and better preserves the original inputs, while generating significantly fewer noise points.
@@ IV-C3 Results on severely occluded objects
To assess our model’s ability to complete objects with severe occlusions, following [16], we remove half of the points that are furthest from the camera in the partial scans and then use these data to train and test our model. As indicated in Table V, the performance of the three self-supervised methods significantly decreases when completing objects with severe occlusions. However, our method still leads the others, achieving the best results in terms of average CD and the CD across six categories. The visual results in Figure 6 demonstrate that our method can still infer the complete shape under extreme occlusion conditions. For example, even when only parts of an airplane’s wing and tail are visible, our method can still reconstruct the fuselage. We believe this is because our model learns category-specific geometric patterns to complete missing parts.
@@ IV-C4 Results on long-tail objects
Our model takes advantage of the common geometries within each category. To test its ability to reconstruct shapes with rare structures, we select long-tail objects of each category for experiments. Specifically, we use a pre-trained point cloud reconstruction model to generate a global feature for each object, and then calculate the cosine similarity between each global shape and others within the same category. The ten samples with the lowest average similarity are selected as the long-tail set for that category, while the ten with the highest average similarity form the common set, representing objects with the most typical geometric structures. As Table VI demonstrates, the performance of the three self-supervised methods drops significantly on the long-tail set, as they rely on common geometric patterns to infer missing structures without supervision from complete point clouds. In addition, our method demonstrates stable performance on the cabinet and couch categories, while on the table category, the performance difference between the two sets is the largest, reaching nearly 34.
@@ IV-C5 Results on novel categories
To evaluate the generalization ability of our model, we train it on all eight categories and then test it on novel categories. Specifically, the testing data are derived from ShapeNet-34 [16], which comprises 34 categories from ShapeNet. As shown in Table VII, the self-supervised methods generally perform poorly on the unseen dataset, indicating their limitations in generalization. This is because, without complete supervision, the self-supervised approaches can only capture some typical geometric structures within the training categories. Even so, our method outperforms ACL-SPC and P2C and demonstrates particularly good performance on the ’keyboard’ category.
@@ IV-C7 Failure cases
In Figure 7, we present some examples of failure cases. When the shape of the target object is rare within its category or the occlusion is severe, our method struggles to reconstruct these unique structures, instead erroneously generating more common structures. This is because, without supervision, our method heavily relies on learning the common geometric features within the category to reconstruct missing structures, which is a constraint shared by current self-supervised methods.
@@ IV-D1 Effects of each module
TABLE IX: Quantitative comparison of the effectiveness of each module in our framework. “✓” and “✕” represent the model with or without the component, respectively. “PE” and “CE” denote the position encoding and the curvature encoding, respectively. “MAN” means the multi-view adversarial network, and “DARE” represents the density-aware radius estimation algorithm. The bold numbers represent the best results.
Component Categories
@@
To analyze the effect of each innovative component in our framework, including position encoding, curvature encoding, multi-view adversarial network, and density-aware radius estimation strategy, we conduct ablation experiments on synthetic data. As Table IX shows, the inclusion of all modules (PE, CE, MAN, and DARE) yields the best average performance, achieving the lowest average error of 9.08. In particular, adding only PE results in a noticeable reduction in the average error to 10.76, especially a substantial gain in the lamp category (from 24.22 to 16.89). We also compared the visualization results with and without using PE to observe the impact of PE on the reconstruction process. As Figure 8 illustrates, without PE, the model predicts some inaccurate points, including outliers that significantly deviate from the partial input and the ground truth. However, after incorporating PE, these inaccurate points are noticeably reduced. We attribute this effect to the following reasons: PE captures the local patterns of the coarse shape $P_{c}$ and the partial input $P_{in}$. As discussed in Section III-A, the PRN retrieves the most relevant PE from $P_{in}$ for $P_{c}$ and uses them to predict an offset for $P_{c}$. Consequently, even though the local regions in $P_{c}$ may be noisy, the PRN leverages the structured and noise-free local patterns from $P_{in}$ to calibrate the positions of noisy local points in $P_{c}$.
@@
Figure 10: Comparison of rendered depth maps with a fixed radius and with the DARE algorithm. The first and second rows show the rendering results before and after applying the DARE algorithm, respectively. The DARE algorithm adaptively adjusts the rendering radius based on the density of projected points, enabling smoother foregrounds for low-density depth maps while preventing the expansion of foreground regions in high-density depth maps. (a)$\sim$(d) High-density depth map. (e)$\sim$(h) Low-density depth map.
Additionally, it is evident from Table IX that incorporating both PE and CE reduces the average CD further to 9.87. CE complements PE by providing additional curvature information, significantly improving performance in categories such as table, couch, and cabinet. Moreover, as observed from the visual results in Figure 9, embedding CE results in a noticeably smoother reconstructed surface when the incomplete input point cloud surface is compact and smooth. Like PE, CE encodes the curvature patterns of the local surfaces of $P_{c}$ and $P_{in}$. Through attention-based retrieval, PRN leverages the most relevant smooth surface knowledge from $P_{in}$ to refine the rough local surfaces of $P_{c}$.
@@
To more intuitively demonstrate the effect of the DARE algorithm, we compared the depth maps rendered with a fixed radius and with the DARE algorithm. As shown in Figure 10, when a fixed radius is used for rendering, low-density depth maps appear smoother, but high-density depth maps often contain many holes. However, after applying the DARE algorithm, the foreground of high-density depth maps becomes smoother, while the edges of foreground pixels in low-density depth maps do not expand. This proves that the proposed DARE algorithm can adaptively select an appropriate point radius based on the density of projected points, thereby rendering higher-quality depth maps.
@@ IV-D2 Ablation on supervisory signals
Our solution employs two types of direct supervisory signals: an incomplete 3D point cloud and a 2D depth map. Although these signals have different data representations, they both originate from the same source, i.e., the incomplete point cloud is derived by back-projecting the depth map into 3D space. We do not incorporate 2D supervision from other viewpoints.
To evaluate the impact of these two supervisory signals on the model’s performance, we conduct ablation studies on the synthetic dataset. We use our complete model and perform experiments by separately removing $L_{part}$ and $L_{rend}$. The results show that non-converge occurs in the absence of either $L_{part}$ or $L_{rend}$. For the case of lacking $L_{rend}$, the CD values achieved on the cabinet, car, couch, and table categories are 36.56, 18.19, 19.68, and 74.83, respectively, while convergence is not achieved on other categories. On the other hand, when $L_{part}$ is removed, CD values on cabinet, chair, car, couch, table, and watercraft categories are 29.13, 21.12, 7.36, 12.00, 12.53, and 9.78, respectively, while convergence failed for plane and lamp categories. We believe that the reason for the non-convergence is twofold. First, the training of GAN itself is inherently unstable [64, 71]. Second, the inclusion of DARE makes training more challenging. This is because DARE requires a relatively accurate initial depth map rendered from the predicted results before estimating a more precise point radius. If the initial depth map has significant deviations, the estimated point radius will also deviate accordingly, leading to significant errors in the second-round rendering. Consequently, the GAN module might collapse, ultimately preventing the model from converging. The lack of either $L_{rend}$ or $L_{part}$, which provides direct supervision, increases the difficulty for the model to predict accurate shapes, contributing to the non-convergence. However, based on the results from the converged categories, $L_{rend}$ exhibits stronger supervisory capabilities compared to $L_{part}$.

### v3

@@ IV-C1 Results on synthetic data
Datasets Methods Average Plane Cabinet Car Chair Lamp Couch Table Watercraft
@@
The quantitative results presented in Table I suggest that our method achieves an impressive average CD of 9.08, which outperforms self-supervised methods and two unpaired approaches (i.e., PCL2PCL and USSPA) by a large margin, and even demonstrates competitive results compared to some fully supervised methods. Regarding the category-specific results, our method significantly outperforms two self-supervised methods across five categories. Specifically, the CD value of MAL-UPC under the lamp category surpasses the results of P2C and ACL-SPC by nearly 5 and 20, respectively. Besides, the CD values of MAL-UPC are lower than those of self-supervised benchmarks by 2-4 under cabinet, couch, and table categories. Furthermore, even with the prior knowledge of complete 3D point clouds, the results of the two unpaired methods, PCL2PCL and USSPA, lag behind our method across multiple categories. In addition, compared to the supervised method FoldingNet, our model achieves a CD difference of less than 3.
@@
Figure 4: Qualitative results on the synthetic dataset. (a) Partial points. (b) Ground truth. (c) FoldingNet [41]. (d) PCN [15]. (e) AnchorFormer [17]. (f) KTNet [29]. (g) USSPA [30]. (h) ACL-SPC [66]. (i) P2C [67]. (j) Ours.
@@
As shown in the last two rows of Table I, the results of our method trained separately for each category are generally better than training them together. However, the latter also outperforms P2C and ACL-SPC. Table II shows that our method achieves the best Precision across most categories, and outperforms self-supervised methods in terms of average Precision. Although it falls slightly behind ACL-SPC in Coverage, the difference remains minimal. Specifically, our Coverage is slightly higher than ACL-SPC by 0.93 (51.38%), while ACL-SPC’s Precision is higher than ours by 7.1 (112.16%). In addition, relying only on Coverage or Precision is insufficient to comprehensively measure the quality of predicted results. For example, if the predicted points fill the whole 3D space, the Coverage value would be zero, and if all the predicted points are concentrated at the position of a single point in the ground truth, the Precision value would be zero. In both cases, the predicted results would be extremely poor. Thus, a comprehensive quantitative evaluation requires combining Coverage and Precision. Therefore, we believe that the slight disadvantage of Coverage does not undermine the overall superiority of our method.
Table III indicates that our method leads on the PCN and CRN results, with the average CD surpassing the second place by 6.04 and 6.09, respectively. However, the performance of our method and ACL-SPC shows a significant decline on the 3D-EPN dataset. To simulate the human perspective, we set the elevation angle for scanning objects between -30 and 30 degrees when generating our dataset, whereas the partial point clouds in 3D-EPN are scanned from arbitrary angles. This difference in viewpoints introduces a domain gap between the datasets, resulting in the underperformance of our method and ACL-SPC. Note that the performance of P2C increases on the 3D-EPN dataset, indicating that it is more robust to viewpoint variations of partial shape scan.
Moreover, as shown in Table I, our method achieves favorable results in the unpaired case, with the average CD significantly outperforming PCL2PCL and USSPA. However, the results slightly lag behind Ours-S model. We believe that, on the one hand, our PRN effectively learns the common 3D structural features within each object and category, thereby reducing the impact of the unpaired complete point cloud. On the other hand, unlike unpaired methods, our MAN leverages a 2D network as the discriminator. It is well-known that 2D neural networks generally possess stronger learning capabilities than 3D MLPs, and our method is able to effectively utilize this advantage.
Furthermore, Table I shows that after adding our rendering loss and MAN module to the P2C model, the average CD decreases by 1.2 and the CD values across six categories also decline. This demonstrates the versatility of our depth map-based strategies, which is also a significant contribution of our work. On the other hand, the average CD of this method is still higher than that of our method by 2.0, further illustrating the superiority of our complete framework over the self-supervised methods.
Figure 4 illustrates a comparison of the visualization results between our method and other benchmarks. We can see that MAL-UPC can infer a more reasonable global shape while better preserving the incomplete parts. For example, the four legs of the input chair point cloud are all thin cylindrical shapes, while self-supervised or even supervised methods reconstruct chair legs as thick rectangular prisms, whereas MAL-UPC preserves its original shape well. In addition, it is obvious that the noise points generated by our method are significantly fewer than those produced by other methods. Self-supervised benchmarks and most supervised approaches generate a large number of outlier points outside the edges of objects, whereas our method has almost no outlier points.
@@ IV-C2 Results on real-world scans
In order to validate the robustness and effectiveness of our method in real-world environments, we evaluate its performance on two challenging datasets: ScanNet and KITTI. As demonstrated in Table IV, compared to ACL-SPC and P2C, our MAL-UPC yields the best UHD on all three categories and the best UCD on the car and table categories. Our method only slightly falls behind the second-best approach in terms of UCD on the chair category. Moreover, regarding the results on the car category, MAL-UPC even surpasses the unpaired and supervised methods by a large margin. Figure 5 shows the qualitative results of our model and other state-of-the-art benchmarks. Similar to its superior performance on synthetic data, our method predicts more reasonable shapes and better preserves the original inputs, while generating significantly fewer noise points.
@@ IV-C3 Results on severely occluded objects
To assess our model’s ability to complete objects with severe occlusions, following [16], we remove half of the points that are furthest from the camera in the partial scans and then use these data to train and test our model. As indicated in Table V, the performance of the three methods significantly decreases when completing objects with severe occlusions. However, our method still leads the others, achieving the best results in terms of average CD and the CD across six categories. The visual results in Figure 6 demonstrate that our method can still infer the complete shape under extreme occlusion conditions. For example, even when only parts of an airplane’s wing and tail are visible, our method can still reconstruct the fuselage. We believe this is because our model learns category-specific geometric patterns to complete missing parts.
@@ IV-C4 Results on long-tail objects
Our model takes advantage of the common geometries within each category. To test its ability to reconstruct shapes with rare structures, we select long-tail objects of each category for experiments. Specifically, we use a pre-trained point cloud reconstruction model to generate a global feature for each object, and then calculate the cosine similarity between each global feature and others within the same category. The ten samples with the lowest average similarity are selected as the long-tail set for that category, while the ten with the highest average similarity form the common set, representing objects with the most typical geometric structures. As Table VI demonstrates, the performance of the three methods drops significantly on the long-tail set, as they rely on common geometric patterns to infer missing structures without supervision from complete point clouds. In addition, our method demonstrates stable performance on the cabinet and couch categories, while on the table category, the performance difference between the two sets is the largest, reaching nearly 34.
@@ IV-C5 Results on novel categories
To evaluate the generalization ability of our model, we train it on all eight categories and then test it on novel categories. Specifically, the testing data are derived from ShapeNet-34 [16], which comprises 34 categories from ShapeNet. As shown in Table VII, the unsupervised methods generally perform poorly on the unseen dataset, indicating their limitations in generalization. This is because, without complete shapes as supervision, these approaches can only capture some typical geometric structures within the training categories. Even so, our method outperforms ACL-SPC and P2C and demonstrates particularly good performance on the keyboard category.
@@ IV-C7 Failure cases
In Figure 7, we present some examples of failure cases. When the shape of the target object is rare within its category or the occlusion is severe, our method struggles to reconstruct the accurate structures, instead erroneously generating more common structures. This is because, without full supervision, our method heavily relies on learning the common geometric features within the category to reconstruct missing structures, which is a constraint shared by current self-supervised methods.
@@ IV-D1 Effects of each module
TABLE IX: Quantitative comparison of the effectiveness of each module in our framework. “✓” and “✗” represent the model with or without the component, respectively. “PE” and “CE” denote the position encoding and the curvature encoding, respectively. “MAN” means the multi-view adversarial network, and “DARE” represents the density-aware radius estimation algorithm. The bold numbers represent the best results.
Components Categories
@@
To analyze the effect of each innovative component in our framework, including position encoding, curvature encoding, multi-view adversarial network, and density-aware radius estimation algorithm, we conduct ablation experiments on synthetic data. As Table IX shows, the inclusion of all modules (PE, CE, MAN, and DARE) yields the best average performance, achieving the lowest average CD of 9.08. In particular, adding only PE results in a noticeable reduction in the average CD to 10.76, especially a substantial gain in the lamp category (from 24.22 to 16.89). We also compare the visualization results with and without using PE to observe the impact of PE. As Figure 8 illustrates, without PE, the model predicts some inaccurate points, including outliers that significantly deviate from the partial input and the ground truth. After incorporating PE, however, these inaccurate points are noticeably reduced. We attribute this effect to the following reasons: PE captures the local patterns of the coarse shape $\textbf{P}_{c}$ and the partial input $\textbf{P}_{in}$. As discussed in Section III-A, the PRN retrieves the most relevant PE from $\textbf{P}_{in}$ for $\textbf{P}_{c}$ and uses them to predict an offset for $\textbf{P}_{c}$. Consequently, even though the local regions in $\textbf{P}_{c}$ may be noisy, the PRN leverages the structured and noise-free local patterns from $\textbf{P}_{in}$ to calibrate the positions of noisy local points in $\textbf{P}_{c}$.
@@
Figure 10: Comparison of rendered depth maps with a fixed radius and with the DARE algorithm. The first and second rows show the rendering results before and after applying the DARE algorithm, respectively. The DARE algorithm adaptively adjusts the rendering radius based on the density of projected points, enabling smoother foregrounds for low-density depth maps while preventing the expansion of foreground regions in high-density depth maps. (a)$\sim$(d) High-density depth maps. (e)$\sim$(h) Low-density depth maps.
Additionally, it is evident from Table IX that incorporating both PE and CE reduces the average CD further to 9.87. CE complements PE by providing additional curvature information, significantly improving performance in categories such as table, couch, and cabinet. Moreover, as observed from the visual results in Figure 9, embedding CE results in a noticeably smoother reconstructed surface when the incomplete input point cloud surface is compact and smooth. Like PE, CE encodes the curvature patterns of the local surfaces of $\textbf{P}_{c}$ and $\textbf{P}_{in}$. Through attention-based retrieval, PRN leverages the most relevant smooth surface knowledge from $\textbf{P}_{in}$ to refine the rough local surfaces of $\textbf{P}_{c}$.
@@
To more intuitively demonstrate the effect of the DARE algorithm, we compare the depth maps rendered with a fixed radius and with the DARE algorithm. As shown in Figure 10, when a fixed radius is used for rendering, high-density depth maps appear smoother, but low-density depth maps often contain many holes. However, after applying the DARE algorithm, the foreground of low-density depth maps becomes smoother, while the edges of foreground pixels in high-density depth maps do not expand. This proves that the proposed DARE algorithm can adaptively select an appropriate point radius based on the density of projected points, thereby rendering higher-quality depth maps.
@@ IV-D2 Ablation on supervisory signals
Our solution employs two types of direct supervisory signals: an incomplete 3D point cloud and a 2D silhouette map. Although these signals have different data representations, they both originate from the same source, i.e., the incomplete point cloud is derived by back-projecting the depth map into 3D space, and the silhouette map is binarized from the depth map. We do not incorporate 2D supervision from other viewpoints.
To evaluate the impact of these two supervisory signals on the model’s performance, we conduct ablation studies on the synthetic dataset. We use our complete model and perform experiments by separately removing $L_{part}$ and $L_{rend}$. The results show that non-converge occurs in the absence of either $L_{part}$ or $L_{rend}$. For the case of lacking $L_{rend}$, the CD values achieved on the cabinet, car, couch, and table categories are 36.56, 18.19, 19.68, and 74.83, respectively, while convergence is not achieved on other categories. On the other hand, when $L_{part}$ is removed, the CD values on cabinet, chair, car, couch, table, and watercraft categories are 29.13, 21.12, 7.36, 12.00, 12.53, and 9.78, respectively, while convergence failed for plane and lamp categories. We believe that the reasons for the non-convergence are twofold. First, the training of GAN is inherently unstable [64, 71]. Second, the inclusion of DARE makes training more challenging. This is because DARE requires a relatively accurate initial depth map rendered from the predicted results before estimating a more precise point radius. If the initial depth map has significant deviations, the estimated point radius will also deviate accordingly, leading to significant errors in the second-round rendering. Consequently, the GAN module might collapse, ultimately preventing the model from converging. In conclusion, both $L_{rend}$ and $L_{part}$ are required to predict accurate shapes. However, based on the results from the converged categories, $L_{rend}$ exhibits stronger supervisory capabilities compared to $L_{part}$.